OpenStudy (anonymous):

What is the vertex of the parabola defined by the equation y = x2 + 4x + 4?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

first coordinate of the vertex is \(-\frac{b}{32a}=-\frac{4}{2\times 1}=-2\)

OpenStudy (anonymous):

second coordinate of the vertex is what you get when you replace \(x\) by \(-2\) and compute \[(-2)^2+4\times (-2)+4\]